### v1

# Quantifier Elimination and Invariant Theory: Applications to Quaternions, Octonions, and Other Algebras

Abstract
We build on our previous paper [constructive] by using the general method introduced there in conjunction with invariant theory. This yields quantifier elimination results for the classical quaternions, octonions, as well as other classes of finite-dimensional algebras over real closed and algebraically closed fields. In particular, the first two examples answer an open question posed recently in [savi].
1 Introduction and Preliminaries
1.1 Introduction
Quantifier elimination is a central topic in model theory, with deep connections to algebra, geometry, and computation, especially in the real case [basu, boch, pd]. Classical results establish quantifier elimination for algebraically closed fields and real closed fields in the languages $\mathcal{F}:=(+,-,\times,0,1)$ and $(\mathcal{F},\leq)$ respectively.
In contrast, much less is known about quantifier elimination for noncommutative algebras. Even for natural and well-studied structures such as the classical quaternions $\mathbb{H}(k)$ or octonions $\mathbb{O}(k)$, for $k$ real closed, the model-theoretic behavior is considerably more subtle. In particular, it was recently shown in [savi] that their first-order theories in the pure ring language do not admit quantifier elimination, and the problem of identifying suitable language extensions was left open.
In this paper, we develop a systematic method for obtaining quantifier elimination for finite-dimensional algebras via invariant theory. Building on the constructive framework introduced in [constructive], we show that the existence of sufficiently expressive invariants for diagonal automorphism actions allows one to transfer quantifier elimination from the base field to the algebra.
As applications, we prove quantifier elimination for the classical quaternions and octonions over real closed fields in a natural language expansion by conjugation $\overline{}$ and order $\leq$. More generally, we obtain criteria for finite-dimensional algebras $\mathcal{A}$ over real closed and algebraically closed fields, connecting properties of their invariant algebras $k[\mathcal{A}^{m}]^{G}$ for $G<\text{Aut}(\mathcal{A})$, to possible candidates of language extensions of $\mathcal{F}$, such that quantifier elimination holds.
The method is constructive whenever quantifier elimination in the base field is constructive and our "transfer function" (see Theorem˜2.1) is explicitly known. In particular, for the case of real quaternions, the proof naturally provides an algorithm that converts the problem in $\mathbb{H}(k)$ of eliminating $l$ quantified variables from the conjunction of $r$ atomic formulas in $l+m$ variables, to quantifier elimination of $4(l+m)$ variables from the conjunction of $4r+m+{m^{2}}+{m^{3}}$ formulas in $4(m+l)+m+{m^{2}}+{m^{3}}$ variables in the base field $k$. For octonions $\mathbb{O}(k)$ the same holds true, by replacing $4$ by $8$ and $m+{m^{2}}+{m^{3}}$ by $o_{m}=\mathcal{O}(m^{4})$. For nice structure on the formula (for example $\exists$-block structure), the complexity of the corresponding algorithm should therefore be realistic for implementation (compare also to the complexity analysis of [constructive]).
1.2 Preliminaries from Algebra
We briefly introduce real closed fields, algebraically closed fields, quaternions and octonions.
A field $k$ is algebraically closed if every nonconstant polynomial over $k$ has a zero in $k$. The most prominent example is $k=\mathbb{C}$, but every field is contained in an, up to isomorphism, unique algebraically closed field, denoted its algebraic closure $\bar{k}$.
A field $k$ is real closed if 1-1 is not a square in $k$ and the field extension $k(\sqrt{-1})$ is algebraically closed (in which case it is the algebraic closure of $k$). There are other characterizations [pd], but this one is the easiest to state. Here, $\mathbb{R}$ is the most prominent example. Real closed fields have a unique field ordering, where an element is nonnegative if and only if it is a square. Every ordered field is contained in a real closed field.
If $k$ is real closed, then on $k(\sqrt{-1})$ one obtains an involution, just as complex conjugation on $\mathbb{C}$, by multiplying the coefficient of $\sqrt{-1}$ by 1-1. As we will see below, this also applies to the quaternions $\mathbb{H}$ and octonions $\mathbb{O}$ over $k$.
For the remainder of this subsection, $k$ denotes a real closed field.
The (non-split, classical) quaternions over $k$, denoted $\mathbb{H}(k)$, form the $k$-algebra generated by elements $\mathbf{i},\mathbf{j}$ satisfying
$\mathbf{i}^{2}=\mathbf{j}^{2}=-1,\qquad\mathbf{i}\mathbf{j}=-\mathbf{j}\mathbf{i}.$
Setting $\mathbf{k}=\mathbf{i}\mathbf{j}$, one obtains
$\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=-1,\qquad\mathbf{i}\mathbf{j}=\mathbf{k}=-\mathbf{j}\mathbf{i}.$
Every element of $\mathbb{H}(k)$ has a unique expression
$x=a+b\mathbf{i}+c\mathbf{j}+d\mathbf{k},\qquad a,b,c,d\in k,$
i.e. the underlying vector space is $k^{4}$.
In this vector space representation we introduce the notation $\operatorname{Re}(x):=a$ and $\operatorname{Im}(x):=(b,c,d)$.
$\mathbb{H}(k)$ is central simple and associative over $k$. It carries a standard involution (conjugation)
$\overline{x}=a-b\mathbf{i}-c\mathbf{j}-d\mathbf{k},$
whose associated quadratic form
$N(x)=x\overline{x}=a^{2}+b^{2}+c^{2}+d^{2}$
is positive definite and multiplicative. In particular, $\mathbb{H}(k)$ is, up to isomorphism, the unique $4$-dimensional division algebra over $k$ and a composition algebra.
The (classical, non-split) octonions over $k$, denoted $\mathbb{O}(k)$, are obtained from $\mathbb{H}(k)$ by the Cayley–Dickson doubling process. As a $k$-vector space,
$\mathbb{O}(k)=\mathbb{H}(k)\oplus\mathbb{H}(k)\boldsymbol{\ell},\qquad\boldsymbol{\ell}^{2}=-1,$
with multiplication defined by
$(a+b\boldsymbol{\ell})(c+d\boldsymbol{\ell})=(ac-\overline{d}b)+(da+b\overline{c})\boldsymbol{\ell},\qquad a,b,c,d\in\mathbb{H}(k).$
The algebra $\mathbb{O}(k)$ is $8$-dimensional, noncommutative, nonassociative but alternative. Just as for quaternions, we use the vector space structure to define, for $x=a+b\boldsymbol{\ell}$,
$\operatorname{Re}(x):=\operatorname{Re}(a),\qquad\operatorname{Im}(x):=(\operatorname{Im}(a),\operatorname{Re}(b),\operatorname{Im}(b)).$
In some literature (for example [schwarzoct, alginv]), $2\operatorname{Re}$ is also denoted as the trace $\operatorname{tr}$.
Conjugation extends by
$\overline{a+b\boldsymbol{\ell}}=\overline{a}-b\boldsymbol{\ell},$
and the norm
$N(x)=x\overline{x}$
remains multiplicative and positive definite. Thus $\mathbb{O}(k)$ is a composition algebra and is the unique $8$-dimensional division algebra over $k$.
1.3 Preliminaries from Logic
The model theory, used throughout this paper will be very light. We will work in the classical setup of first-order logic and model theory, as described for example in [cha, hod, pre]. We denote formal languages by calligraphic letters such as $\mathcal{L}$.
A language can be extended by an additional symbol $S$; we denote this by $(\mathcal{L},S)$. In order to speak of rings, a language $\mathcal{L}$ needs at least two function symbols $\cdot$ of arity $2$, and two constant symbols $0,1$. With these symbols one can write down the ring axioms and the axioms of algebraically closed or real closed fields, for example. Throughout this paper, we denote this specific language by $\mathcal{F}$. The complex numbers $\mathbb{C}$ form an obvious $\mathcal{F}$-structure, with the usual interpretation of $\cdot,0,1$.
If we want to talk about ordered rings, we can extend the language by the relation symbol $\leq$. An obvious $(\mathcal{F},\leq)$-structure is $\mathbb{R}$ with the usual interpretations. Note that for real closed fields, the extension by $\leq$ is to some extent redundant, since nonnegative elements are squares and nonnegativity can thus be expressed in $\mathcal{F}$ alone. However, this uses a quantifier, and to obtain quantifier elimination one needs to include the relation symbol $\leq$.
To denote a formula we will use small greek letters like $\phi$ and to emphasize the free variables $x_{1},\dots,x_{m}$ of a formula $\phi$, we sometimes write $\phi(x_{1},\dots,x_{m})$. Given such a formula and a structure $A$, $\phi$ defines the subset of $A^{m}$ of all $m$-tuples that satisfy the formula, and we write $\phi(A)$ for this subset. Such sets are called definable.
We say an $\mathcal{L}$-structure $A$ admits quantifier elimination if every definable set is also definable without quantifiers. More precisely, for every $\mathcal{L}$-formula $\phi(x_{1},\dots,x_{m})$ there exists a quantifier-free $\mathcal{L}$-formula $\psi(x_{1},\dots,x_{m})$ with
$\phi(A)=\psi(A).$
It is well known that algebraically closed fields, i.e. $\mathcal{F}$-structures satisfying the axioms of algebraically closed fields, and real closed fields, i.e. $(\mathcal{F},\leq)$-structures satisfying the axioms of real closed fields, admit quantifier elimination. In fact, a stronger result holds: quantifiers can be removed independently of the actual field (that is, the theory of algebraically closed or real closed fields admits quantifier elimination).
Both results can be proven by showing that the projection of a definable set is again definable. For algebraically closed fields, definable sets are constructible sets in the Zariski topology, while for real closed fields they are the semialgebraic sets. Since projections are defined by formulas of the form $\exists x\,\phi(x,x_{1},\dots,x_{m})$, the general case can then be deduced inductively. In the case of algebraically closed fields, a constructive proof reduces to a simple application of the Euclidean algorithm. For real closed fields the result is less trivial and is the content of the Tarski–Seidenberg Theorem. Also in this case constructive proofs exist (see for example [basu, boch, sch]).
Let $k$ be a field in the language $\mathcal{F}$. Any unital $k$-algebra can also be considered as a structure over the language $\mathcal{F}$ by interpreting the symbols $+,-,\times$ as the respective algebra addition, subtraction, and multiplication, and $1$ and 0 as $1_{\mathcal{A}}$ and $0_{\mathcal{A}}$, respectively.
If $k$ is an ordered field in the language $(\mathcal{F},\leq)$, a finite-dimensional $k$-algebra can also be defined over this extended language by interpreting $\leq$ as the order naturally defined on $k1_{\mathcal{A}}$. For the base field $k$ in general one can only single out the constants $c1_{\mathcal{A}}$ for $c\in\mathbb{Q}$ or $\mathbb{F}_{p}$, depending on the characteristic of $k$. This is why one sometimes sees the language extension $(\mathcal{F},(\lambda_{c})_{c\in k})$, where $\lambda_{c}$ denotes scalar multiplication with $c$. Since our interest is not in axiomatizing the $k$-algebra models, but in interpreting fixed $k$-algebras with a minimal language, allowing for meaningful formulas, we will forego these additional symbols.
If the center
$C(\mathcal{A})=\{a\in\mathcal{A}:\forall x\in\mathcal{A},\ ax=xa\}$
equals $k1_{\mathcal{A}}$ we have an easy way to identify the base field $k$ as a definable subset of $\mathcal{A}$. However, even if $k1_{\mathcal{A}}$ is a strict subset of the center, we can (and will) use the natural isomorphism $\iota:k\to k\cdot 1_{\mathcal{A}}$, for example to transform functions $f:\mathcal{A}^{m}\to k^{n}$ into functions in the $\mathcal{F}$-structure $\mathcal{A}$.
The algebras we will be most interested in are $\mathbb{H}(k)$ and $\mathbb{O}(k)$ for real closed $k$. In the recent paper [savi], general model-theoretic properties have been investigated for theories in the language $\mathcal{F}$ whose models are all isomorphic to $\mathbb{H}(k)$ or $\mathbb{O}(k)$, respectively. In particular, it was shown that these theories do not have quantifier elimination in $\mathcal{F}$, and the question for possible language extensions $\mathcal{F}^{\prime}$ such that quantifier elimination holds, was posed.
We will interpret $\mathbb{H}(k)$ and $\mathbb{O}(k)$ as $(\mathcal{F},\leq,\bar{\ })$-structures, where $\leq$ is interpreted as described above for general $k$-algebras and $\bar{\ }$ as the conjugation, defined in our introductory section on quaternions and octonions. As it will turn out, $(\mathcal{F},\leq,\bar{\ })$ is a language extension allowing for quantifier elimination in both cases.
Just as in our previous paper (although less explicitly stated there), the most important model-theoretic concept for proving this will be that of types and the separation thereof.
Let $A$ be a structure in the language $\mathcal{L}$. We say $\underline{a},\underline{b}\in A^{m}$ have the same $\emptyset$-type if for every $\mathcal{L}$-formula $\phi(x_{1},\dots,x_{m})$ one has
$\underline{a}\in\phi(A)\iff\underline{b}\in\phi(A)$
For a $\mathcal{L}$-structure $A$, a bijective function $f:A\to A$ is called an automorphism if it preserves the respective interpretations of every constant-/function-/relation-symbol. The set of automorphisms forms a group with regards to composition, denoted $\mathrm{Aut}(A)_{\mathcal{L}}$. We will omit the subscript if the language $\mathcal{L}$ is clear from the context.
It is easy to see that the diagonal action of automorphisms preserves types.
Theorem 1.1.
Let $A$ be a structure in the language $\mathcal{L}$ and denote by $\mathrm{Aut}(A)$ its automorphism group. Then for $m$-tuples $\underline{a},\underline{b}\in A^{m}$ one has
$\underline{a}\in\mathrm{Orb}(\underline{b},\mathrm{Aut}(A))\quad\Rightarrow\quad\underline{a}\text{ and }\underline{b}\text{ are of the same }\emptyset\text{-type.}$
1.4 Linear Algebraic Groups and Invariant Theory
We briefly review the notions and results from invariant theory and the theory of linear algebraic groups needed in the sequel. A complete treatment is impossible here, the references we used include [GIT, bredon, borel, lie, weyl, grot, sym]. The modern treatment of geometric invariant theory is heavily scheme-theoretic but for the purpose of our paper it is luckily possible to avoid the more technical details. We will state all the relevant results, such that they are readable without any background in algebraic geometry. In that spirit we start with the, not maximally general, definition of a linear algebraic group.
Definition 1.1.
Let $k$ be an algebraically closed field of characteristic 0 and $G<\mathrm{GL}_{n}(k)$. We say that $G$ is a linear algebraic group (over $k$) if it is Zariski closed in $\mathrm{GL}_{n}(k)$, i.e. if there exists an ideal $I$ of polynomial functions in the matrix entries such that
$G=\mathcal{V}(I)\subset\mathrm{GL}_{n}(k).$
The coordinate ring of $G$ is defined as
$k[G]=k[\mathrm{GL}_{n}(k)]/\sqrt{I},$
where
$k[\mathrm{GL}_{n}(k)]=k[(M_{ij})_{i,j\leq n},\text{det}^{-1}]$
and $\sqrt{I}=\mathcal{I}(G)$ denotes the ideal of polynomial functions vanishing on $G$.
If $l$ is a subfield of $k$ and the ideal $\mathcal{I}(G)$ is already generated (as $k$-ideal) by the corresponding ideal $\mathcal{I}(G)_{l}$, we say that $G$ is defined over $l$. In this case we can define the group of $l$-rational points
$G_{l}=G\cap\mathrm{GL}_{n}(l).$
Conversely if $G$ is a linear algebraic group defined over a field $k$, one can for field extension $K$ of $k$ also define define the $K$-rational points $G_{K}$, by base change. This is especially relevant, when allowing for linear algebraic groups over arbitrary fields.
A more general notion is that of an $l$-form of $G$ (for details see [borel] or [sym] for the case $l=\mathbb{R}$).
As an example, the group $\mathrm{GL}_{n}(\mathbb{C})$ is defined over $\mathbb{R}$, its $\mathbb{R}$-rational points are $\mathrm{GL}_{n}(\mathbb{R})$, yet it has many more $\mathbb{R}$-forms, for example the indefinite unitary groups $\mathrm{U}(p,q)$ for $p+q=n$.
Another example, which will become important later, is the exceptional simple linear algebraic group $G_{2}$. The group $G_{2}(\mathbb{C})$ can be defined as the automorphism group of the complex octonions $\mathbb{O}(\mathbb{C})$. It is defined over $\mathbb{R}$, and the group of $\mathbb{R}$-rational points is $G_{2(2)}$, which is the automorphism group of the real split octonions. There is, however, another $\mathbb{R}$-form, namely the compact Lie group $G_{2}(\mathbb{R})$, which is the automorphism group of the real division octonions $\mathbb{O}(\mathbb{R})$, which we defined in the previous section.
For later results, the notion of a reductive group will play a central role. There are many equivalent characterizations, we again give one for the case where $k$ is algebraically closed.
Definition 1.2.
A linear algebraic group $G$ is called reductive if its largest smooth connected unipotent normal subgroup is trivial.
We now recall the notion of a regular (also called rational) representation.
Definition 1.3.
Let $k$ be an algebraically closed field and $G$ a linear algebraic group. A representation $(V,\pi)$ is called regular if $V$ is a finite-dimensional $k$-vector space and the map
$G\times V\to V,\qquad(g,v)\mapsto\langle v^{*},\pi(g)v\rangle$
is a regular morphism, i.e. belongs to $k[V]$. (Remember, that by finite dimensionality $V\cong V^{*}$).
By definition, obviously, for any linear algebraic group $G<\mathrm{GL}_{n}(k)$ acting naturally on $k^{n}$, the corresponding representation is regular.
Let $(V,\pi)$ be a regular representation of a linear algebraic group $G$. The action of $G$ on $V$ induces a natural action on the coordinate ring $k[V]$ defined by
$(g\cdot f)(v):=f(\pi(g^{-1})v).$
Definition 1.4.
Let $G$ be a linear algebraic group acting by means of a regular representation on a finite-dimensional $k$-vector space $V$. The invariant algebra is the subalgebra
$k[V]^{G}=\{f\in k[V]\mid g\cdot f=f\ \text{for all }g\in G\}.$
The most important example for us arises when $G<\mathrm{Aut}(\mathcal{A})$ for a finite-dimensional $k$-algebra $\mathcal{A}$. Since $\mathrm{Aut}(\mathcal{A})\subset\mathrm{GL}(\mathcal{A})$, it is a linear algebraic group and the natural action, gives a regular representation on $\mathcal{A}$. This action extends diagonally to $\mathcal{A}^{m}$, yielding a regular representation and the corresponding invariant algebra $k[\mathcal{A}^{m}]^{G}$.
Remark 1.1.
It should be noted that linear algebraic groups, reductiveness and regular representations are definitions we introduced, as they are often naturally given in our setting and also requirements for Theorem˜1.3 and Theorem˜1.2. The notion of an invariant algebra can, however, be defined in exactly the same way for a general field $k$ and any group $G$ acting on a finite-dimensional $k$-vector space $V$ and we denote it again with $k[V]^{G}$.
If two points $v,w\in V$ lie in the same $G$-orbit, then clearly $f(v)=f(w)$ for all $f\in k[V]^{G}$. The following fundamental result describes when the converse holds.
Theorem 1.2.
[GIT] Let $k$ be an algebraically closed field, $G$ a reductive linear algebraic group, and $(V,\pi)$ a regular representation. Then the invariant algebra separates orbit closures:
$\forall f\in k[V]^{G}:\ f(v)=f(w)\quad\Longleftrightarrow\quad\overline{\mathrm{Orb}(v,G)}\cap\overline{\mathrm{Orb}(w,G)}\neq\emptyset.$
A second fundamental property of reductive groups acting regularly on $V$ is finite generation of $k[V]^{G}$.
Theorem 1.3.
[sym] Let $k$ be an algebraically closed field, $G$ a reductive linear algebraic group, and $(V,\pi)$ a regular representation. Then the invariant algebra $k[V]^{G}$ is finitely generated.
We now turn to compact real Lie groups acting on real algebras. Real Lie groups can be seen as special cases of linear algebraic groups, when generalizing to linear algebraic groups over arbitrary fields. This generalization is however somewhat more involved and the real versions of Theorem˜1.3 and Theorem˜1.2, that we will show below, are in fact simpler and historically older, dating back to Hilbert. In contrast to the results discussed above, we will state the results for $\mathbb{R}$ instead of general real closed fields, as this is how they can be found in the literature, and in fact may rely on completeness type arguments, not valid for general real closed fields.
Theorem 1.4.
[gol] Let $V$ be a finite-dimensional $\mathbb{R}$-vector space and $G$ a compact Lie group acting linearly on $V$. Then the invariant algebra $\mathbb{R}[V]^{G}$ is finitely generated.
The next theorem, in the form we use it, reduces to an application of the Stone–Weierstrass theorem.
Theorem 1.5.
[bredon, lie] Let $V$ be a finite-dimensional $\mathbb{R}$-vector space and $G$ a compact Lie group acting on $V^{m}$. Let $p_{1},\dots,p_{n}$ be generators of the invariant algebra $\mathbb{R}[V^{m}]^{G}$. Then the orbit space $V^{m}/G$ is locally compact and Hausdorff, and the, by means of $p_{i}$, induced map
$p:V^{m}/G\to\mathbb{R}^{n}$
is an embedding. In particular, $p$ separates the $G$-orbits on $V^{m}$.
The definitions above are only a glimpse of the very rich and old field of classical and geometric invariant theory. We compare this to the notion of complete invariant we have given in our previous paper, which is much more "hands on" and will therefore remain useful for our results but is also devoid of any theory.
Definition 1.5.
Given an equivalence relation $\sim$ on a set $X$, a function $f:X\to Y$ is called a complete invariant if
$\forall a,b\in X:\quad a\sim b\iff f(a)=f(b).$
If $f$ and $Y$ decompose compatibly as direct products, one also speaks of a complete set of invariants.
2 Results
For sake of completeness, we first restate here the general method introduced in [constructive] for transferring quantifier elimination.
Let $\mathcal{L}$ and $\mathcal{G}$ be formal languages and $A$ and $B$ structures over $\mathcal{L}$ and $\mathcal{G}$, respectively. Assume that for every $m\in\mathbb{N}$ there exists a function
$f_{m}:B^{m}\to A^{d(m)}$
with the following properties:
  1. (i)
    For any definable set $D\subseteq B^{m}$, the set $f_{m}(D)$ is definable in $A^{d(m)}$.
  2. (ii)
    For any quantifier-free definable set $E\subseteq A^{d(m)}$, the set $f_{m}^{-1}(E)\subseteq B^{m}$ is quantifier-free definable.
  3. (iii)
    For any $a\in A^{d(m)}$ and any definable set $D\subseteq B^{m}$, we either have $f_{m}^{-1}(a)\subseteq D$ or $f_{m}^{-1}(a)\cap D=\emptyset$.
We call conditions (i) and (ii) constructive if there is an algorithm transforming definitions of sets into definitions of their image/preimage, as required.
Properties (i) and (ii) allow us to transfer definable sets back and forth via $f_{m}$. Property (iii) basically says that $f_{m}$ is "injective up to definability".
Theorem 2.1.
Let $A,B$ and $(f_{m})_{m\in\mathbb{N}}$ be as described above. Then, if $A$ admits quantifier elimination, so does $B$. Furthermore, if (i) and (ii) are constructive and there exists a constructive quantifier-elimination algorithm for $A$, then there is also a constructive algorithm for quantifier elimination in $B$.
Proof.
Let $D\subseteq B^{m}$ be definable. By (i), $f_{m}(D)$ is definable in $A^{d(m)}$. Since $A$ admits quantifier elimination, $f_{m}(D)$ is quantifier-free definable and so by (ii), $f_{m}^{-1}(f_{m}(D))$ is also quantifier-free definable in $B^{m}$. By (iii), we have $f_{m}^{-1}(f_{m}(D))=D$, which proves the first statement. The statement about constructiveness also follows directly from this argument. ∎
Let in the following $k$ always denote a real closed field. Before applying Theorem˜2.1 to $\mathbb{H}(k)$, we need the following lemma, which provides us with a complete invariant for the diagonal action of $SO_{3}(k)\cong\mathrm{Aut}(\mathbb{H}(k))$ on $\mathbb{H}(k)^{m}$.
Lemma 2.1.
Define
$L_{1}:\mathbb{H}(k)^{m}\to k^{m},\qquad\underline{v}\mapsto(\operatorname{Re}(v_{i}))_{i\leq m},$
$L_{2}:\mathbb{H}(k)^{m}\to k^{m^{2}},\qquad\underline{v}\mapsto\bigl(\operatorname{Re}((v_{i}-\overline{v_{i}})(v_{j}-\overline{v_{j}}))\bigr)_{i,j\leq m},$
$L_{3}:\mathbb{H}(k)^{m}\to k^{m^{3}},\qquad\underline{v}\mapsto\bigl(\operatorname{Re}((v_{i}-\overline{v_{i}})(v_{j}-\overline{v_{j}})(v_{k}-\overline{v_{k}}))\bigr)_{i,j,k\leq m}.$
Then the orbits of the diagonal action of $SO_{3}(k)$ on $\mathbb{H}(k)^{m}$ are separated by $L_{1},L_{2},L_{3}$, i.e.
$L_{1}(\underline{v})=L_{1}(\underline{w})\land L_{2}(\underline{v})=L_{2}(\underline{w})\land L_{3}(\underline{v})=L_{3}(\underline{w})$
if and only if
$\exists R\in SO_{3}(k):\quad R\cdot v_{i}=w_{i}\text{ for all }i\leq m.$
Proof.
For simplicity we assume $k=\mathbb{R}$. Since $SO_{3}(k)$ is definable (even in $\mathcal{F}$ without $\leq$), the statement can be expressed in first-order logic, and thus the general case follows by Tarski’s transfer principle.
The action of $SO_{3}(\mathbb{R})$ is defined by
$R\cdot v=\operatorname{Re}(v)+\langle R\operatorname{Im}(v)^{T},(\mathbf{i},\mathbf{j},\mathbf{k})^{T}\rangle.$
It is clear that $R\cdot\overline{v}=\overline{R\cdot v}$ and $\operatorname{Re}(R\cdot v)=\operatorname{Re}(v)$. This makes the $\Leftarrow$” direction immediate.
For the $\Rightarrow$” direction, note that $L_{1}(\underline{v})=L_{1}(\underline{w})$ gives the clearly necessary condition for the real parts to coincide. Thus the problem reduces to finding a rotation $R\in SO_{3}(\mathbb{R})$ simultaneously transforming $\operatorname{Im}(v_{i})$ to $\operatorname{Im}(w_{i})$.
For $L_{2}$ and $L_{3}$ we observe that by elementary quaternion algebra
$\operatorname{Re}((v-\overline{v})(w-\overline{w}))=-4\langle\operatorname{Im}(v)^{T},\operatorname{Im}(w)^{T}\rangle$
and
$\operatorname{Re}((v-\overline{v})(w-\overline{w})(z-\overline{z}))=-8\langle\operatorname{Im}(v)^{T},\operatorname{Im}(w)^{T}\times\operatorname{Im}(z)^{T}\rangle.$
It is well known that if $m$ points $\underline{p}\in\mathbb{R}^{3}$ have the same pairwise angles and lengths as $m$ other points $\underline{q}\in\mathbb{R}^{3}$, then there exists a transformation $R\in O_{3}(\mathbb{R})$ with $Rp_{i}=q_{i}$ for all $i\leq m$.
To ensure that $R$ can be chosen orientation-preserving (i.e. contained in $SO_{3}(\mathbb{R})$), the condition
$\langle p_{i},(p_{j}\times p_{k}\rangle=\langle q_{i},(q_{j}\times q_{k}\rangle$
for any triple $p_{i},p_{j},p_{k}$ in general position is sufficient. In the case where all points lie in a proper subspace, $R$ can also trivially be chosen to be orientation preserving. This completes the proof. ∎
Corollary 2.1.
The $(\mathcal{F},\leq,\overline{\ })$-structure $\mathbb{H}(k)$ admits quantifier elimination.
Proof.
We apply Theorem˜2.1 to the $(\mathcal{F},\leq,\overline{\ })$-structure $\mathbb{H}(k)$ and the $(\mathcal{F},\leq)$-structure $k$, which admits quantifier elimination.
For this define $d(m)=m+m^{2}+m^{3}$ and
$f_{m}:\mathbb{H}(k)^{m}\to k^{d(m)},\qquad\underline{v}\mapsto L_{1}(\underline{v})\oplus L_{2}(\underline{v})\oplus L_{3}(\underline{v}).$
where we define $L_{i}$ as in Lemma˜2.1. We have to verify conditions (i)–(iii).
(i) is clear. Just write both the defining formula and the definition of $f_{m}$ down in the vector entries and use existential quantifiers to obtain the image.
For (ii), suppose $\phi(x_{1},\dots,x_{d(m)})$ is a quantifier-free $(\mathcal{F},\leq)$-formula defining the set $D\subset k^{d(m)}$. Since
$\operatorname{Re}(v)=\iota^{-1}\!\left(\frac{1}{2}(v+\overline{v})\right),$
each coordinate of $\iota^{(m)}(f_{m}(y_{1},...,y_{m}))$ is a simple $(\mathcal{F},\overline{\ })$-term. Therefore
$\phi\bigl(\iota(f_{m}(y_{1},\dots,y_{m})_{1}),...,\iota(f_{m}(y_{1},\dots,y_{m})_{d(m)})\bigr)$
is a quantifier-free $(\mathcal{F},\leq,\overline{\ })$-formula defining $f_{m}^{-1}(D)$.
For (iii), recall that $f_{m}$ is a complete invariant for the diagonal action of $SO_{3}(k)$ on $\mathbb{H}(k)^{m}$ by Lemma˜2.1. Since for any $R\in SO_{3}(k)$ we have $R\cdot\overline{v}=\overline{R\cdot v}$ and $\operatorname{Re}(R\cdot v)=\operatorname{Re}(v)$, it follows that
$\mathrm{Aut}(\mathbb{H}(k))_{(\mathcal{F},\leq,\overline{\ })}=\mathrm{Aut}(\mathbb{H}(k))_{\mathcal{F}}=SO_{3}(k).$
Hence by Theorem˜1.1 the fibers of $f_{m}$ have constant $\emptyset$-type, in particular, they cannot be split by definable sets. ∎
We can repeat the basic proof idea for the octonions $\mathbb{O}(k)$, almost the same. But first, we again need to determine a complete invariant, which turns out a little more complicated than in the quaternion case.
Lemma 2.2.
Define
$L:\mathbb{O}(k)^{m}\to k^{I(m)},\qquad\underline{v}\mapsto\bigl(\operatorname{Re}(\omega_{i}(v_{1},\dots,v_{m}))\bigr)_{i\in I(m)},$
where $I(m)$ denotes the set of all possible (nonassociative) words in $m$ letters of length $\leq 4$.
Then the orbits of the diagonal action of $\mathrm{Aut}(\mathbb{O}(k))$ on $\mathbb{O}(k)^{m}$ are separated by $L$, i.e.
[…]
Proof.
Again we first assume $k=\mathbb{R}$.
We have
$\mathrm{Aut}(\mathbb{O}(\mathbb{R}))=G_{2}(\mathbb{R}),$
which is a real compact Lie group and can be defined as the stabilizer of the action $SO_{7}(\mathbb{R})\curvearrowright\mathbb{O}(\mathbb{R})$ with respect to the $3$-form $\phi=e^{123}+e^{145}+e^{167}+e^{246}-e^{257}-e^{347}-e^{356},$ where $e^{ijk}$ denotes the exterior product of the corresponding dual basis vectors of $\mathbb{R}^{7}$.
Its Lie group complexification is the connected and reductive linear algebraic group
$G_{2}(\mathbb{C})=\mathrm{Aut}(\mathbb{O}(\mathbb{C})).$
$G_{2}(\mathbb{R})$ is a Zariski dense $\mathbb{R}$-form in $G_{2}(\mathbb{C})$ and it naturally acts on $\mathbb{O}(\mathbb{C})$ by $g\otimes\mathrm{id}_{\mathbb{C}}$ (see [sym] for density of real forms or also [borel, grot] for more general statements regarding unirationality).
Now it has been shown in [schwarzoct] (see also [alginv]) that $\mathbb{C}[\mathbb{O}(\mathbb{C})^{m}]^{G_{2}(\mathbb{C})}$ is generated by the forms
$\operatorname{Re}(\omega(x_{1},\dots,x_{m})),$
where $\omega$ ranges over all nonassociative words of length $\leq 4$. Note that all those forms, written as homogeneous polynomials in the vector entries, have real valued coefficients.
Again the $\Leftarrow$” direction is clear, since $G_{2}(\mathbb{R})\curvearrowright\mathbb{O}(\mathbb{R})^{m}$ fixes the real parts.
For the $\Rightarrow$” direction, assume $L(\underline{v})=L(\underline{w})$. Passing to the complexification $\mathbb{O}(\mathbb{C})^{m}=(\mathbb{O}(\mathbb{R})\otimes_{\mathbb{R}}\mathbb{C})^{m}$, it then follows from [schwarzoct] that
$f(v_{1}\otimes 1_{\mathbb{C}},\dots,v_{m}\otimes 1_{\mathbb{C}})=f(w_{1}\otimes 1_{\mathbb{C}},\dots,w_{m}\otimes 1_{\mathbb{C}})\quad\text{for all }f\in\mathbb{C}[\mathbb{O}(\mathbb{C})^{m}]^{G_{2}(\mathbb{C})}.$
But by Zariski density, every $G_{2}(\mathbb{R})$-invariant function in $\mathbb{C}[\mathbb{O}(\mathbb{C})^{m}]$ is also $G_{2}(\mathbb{C})$-invariant (the evaluation function $F_{v}(g):=f(g\cdot v)$ for arbitrary fixed $v\in\mathbb{O}(\mathbb{C})^{m}$ and $f\in\mathbb{C}[\mathbb{O}(\mathbb{C})^{m}]^{G_{2}(\mathbb{C})}$ is clearly regular) and we get
$f(v_{1}\otimes 1_{\mathbb{C}},\dots,v_{m}\otimes 1_{\mathbb{C}})=f(w_{1}\otimes 1_{\mathbb{C}},\dots,w_{m}\otimes 1_{\mathbb{C}})\quad\text{for all }f\in\mathbb{R}[\mathbb{O}(\mathbb{C})^{m}]^{G_{2}(\mathbb{R})}\subset\mathbb{C}[\mathbb{O}(\mathbb{C})^{m}]^{G_{2}(\mathbb{R})}$
So by restricting back to the real subspace $\mathbb{O}(\mathbb{R})\otimes 1_{\mathbb{C}}$, we can use compactness of $G_{2}(\mathbb{R})$ to apply Theorem˜1.5:
$\exists R\in G_{2}(\mathbb{R}):\quad(R\otimes id_{\mathbb{C}})\cdot(v_{i}\otimes 1_{\mathbb{C}})=w_{i}\otimes 1_{\mathbb{C}}\text{ for all }i\leq m.$
and the desired result follows.
Again the result for general real closed fields follows by Tarski’s transfer principle as $G_{2}(k)$ is a definable subset of $k^{49}$, as can be seen for example by its definition as stabilizer subgroup. ∎
Corollary 2.2.
The $(\mathcal{F},\leq,\overline{\ })$-structure $\mathbb{O}(k)$ admits quantifier elimination.
Proof.
We proceed exactly as in the proof of Corollary˜2.1, mutatis mutandis, again using that $\operatorname{Re}(v)=\iota^{-1}\!\left(\frac{1}{2}(v+\overline{v})\right).$
Remark 2.1.
For both quaternions and octonions the symbol $\overline{\ }$ was needed to be able to express the invariant in our extended language. Note that $\overline{\ }$, $\operatorname{Re}$, $\operatorname{Im}$, and $N$ are interdefinable, so the precise choice of the additional symbol can be handled flexibly. On the other hand it was essential for our proof to extend $\mathcal{F}$ by $\leq$ and interpret it as the partial order on the center. This had to be done to ensure condition (ii) of Theorem˜2.1 and indeed, the symbol $\overline{}$ allows for defining a copy of $k$ in the $(\mathcal{F},\overline{})$-structure $\mathbb{H}(k)$ or $\mathbb{O}(k)$ without quantifiers. Thus formulas like
$\exists y:\text{Re}(x)=\text{Re}(y)\text{Re}(y)$
suggest the necessity for the addition of the order symbol $\leq$.
For the special cases of $\mathbb{H}(k)$ and $\mathbb{O}(k)$, we explicitly determined complete invariants in order to state the result for a concrete language extension of $\mathcal{F}$ and also to preserve constructiveness. If we give up on these two features, we can use the general invariant theory developed above to obtain more general statements.
Theorem 2.2.
Let $k$ be a real closed field in the language $(\mathcal{F},\leq)$ and $\mathcal{A}$ a finite-dimensional $k$-algebra. Assume that
$G<\mathrm{Aut}(\mathcal{A})_{(\mathcal{F,\leq})}$
is a $(\mathcal{F},\leq)$-definably compact subset of $k^{n^{2}}$ with respect to the order topology. Let $S_{m}$ be a finite generating set of $k[\mathcal{A}^{m}]^{G}$ for each $m\in\mathbb{N}$. Choose $(\mathcal{F},\leq,E)$ to be an extended language such that the interpretation of each $e\in E$ for $\mathcal{A}$ lies in $\iota(k[\mathcal{A}^{\text{arity}(e)}]^{G})$ and for all $s\in S_{m}$, $\iota(s)$ corresponds to the evaluation of a $(\mathcal{F},\leq,E)$-term. Then the $(\mathcal{F},\leq,E)$-structure $\mathcal{A}$ admits quantifier elimination.
Proof.
Define
$f_{m}:\mathcal{A}^{m}\to k^{|S_{m}|},\qquad f_{m}(\underline{a})_{i}=s_{i}(\underline{a}).$
We want to apply Theorem˜2.1. Since $k$ is real closed, it admits quantifier elimination in $(\mathcal{F},\leq)$, so it only remains to verify (i)–(iii).
Condition (i) holds because $\mathcal{A}$ is a $k$-algebra and the functions interpreted by $E$ are by definition polynomials in the $k$-valued coordinates. Again we use existential quantifiers to obtain the image.
(ii) follows by assumption on the extension $E$. If $\phi(x_{1},\dots,x_{|S_{m}|})$ is a quantifier-free $(\mathcal{F},\leq)$-formula defining $D\subset k^{|S_{m}|}$, then
$\phi\bigl(\iota(s_{1}(y_{1},\dots,y_{m})),\dots,\iota(s_{|S_{m}|}(y_{1},\dots,y_{m}))\bigr)$
is a quantifier-free $(\mathcal{F},\leq,E)$-formula defining $f_{m}^{-1}(D)$.
For (iii), first assume $k=\mathbb{R}$. By [comp], in $o$-minimal structures definable compact sets can equivalently be defined as definable, closed and bounded sets. These are first-order definable properties in $(\mathcal{F},\leq)$ and therefore can be transferred to classical compactness on $\mathbb{R}$ by Heine-Borel.
Then, since $G$ is a compact subgroup of $\mathrm{Aut}(\mathcal{A})<\mathrm{GL}_{n}(k)$ it is a Lie group and acts diagonally on $\mathcal{A}^{m}$. Thus by Theorem˜1.5 $k[\mathcal{A}^{m}]^{G}$ separates the orbits of the diagonal action $G\curvearrowright\mathcal{A}^{m}$. Since $(s_{i})$ generate $k[\mathcal{A}^{m}]^{G}$ we get the following relation:
$f_{m}(\underline{a})=f_{m}(\underline{b})\Leftrightarrow i(\underline{a})=i(\underline{b})\forall i\in k[\mathcal{A}^{m}]^{G}\Leftrightarrow\exists g\in G:g\underline{a}=\underline{b}$
By Theorem˜1.1 and since $G<\text{Aut}(\mathcal{A})_{(\mathcal{F},\leq)}$ this implies that $\underline{a}$ and $\underline{b}$ are of the same $\emptyset$-type in $(\mathcal{F},\leq)$ and since all functions interpreted by $E$ are $G$-invariant $\underline{a}$ and $\underline{b}$ are also of the same $\emptyset$-type in $(\mathcal{F},\leq,E)$. By definability of $G$ the statement for general real closed $k$ follows by Tarski’s transfer principle, which finishes the proof of $(iii)$. ∎
Remark 2.2.
The conditions for the language extension $E$ Theorem˜2.2 seem somewhat far fetched. However, note that so called first fundamental theorems (see [sym]) give explicit descriptions of generators for many invariant algebras of $k$-algebras, and their corresponding automorphism groups (or subgroups thereof) by means of tensor forms. These are usually stated for the algebraically closed case, however, it should be possible to often use Zariski denseness arguments, as we did in the proof of Lemma˜2.2, to infer back to the real closed case.
What about algebras $\mathcal{A}$ over an algebraically closed field (in the following always denoted $k$)? Compactness of the acting group was essential in the real closed case. Comparing with Theorem˜1.2, we see that in general $\mathrm{Aut}(\mathcal{A})$ (or subgroups thereof) separate only orbit closures, which need not separate $\emptyset$-types. Unfortunately compactness for linear algebraic groups $G$ over $k$, only holds in the case of $G$ finite, restricting us quite a bit.
Theorem 2.3.
Let $k$ be an algebraically closed field in the language $\mathcal{F}$ and $\mathcal{A}$ a finite-dimensional $k$-algebra. Assume $G<\mathrm{Aut}(\mathcal{A})_{\mathcal{F}}$ is a finite linear algebraic group. Let $S_{m}$ be a minimal generating set of $k[\mathcal{A}^{m}]^{G}$ for each $m\in\mathbb{N}$, and let $(\mathcal{F},E)$ be an extended language such that the interpretation of each $e\in E$ in $\mathcal{A}$ lies in $\iota(k[\mathcal{A}^{\text{arity}(e)}]^{G})$ and for all $s\in S_{m}$, $\iota(s)$ corresponds to evaluation of a $(\mathcal{F},E)$-term. Then the $(\mathcal{F},E)$-structure $\mathcal{A}$ admits quantifier elimination.
Proof.
First note that by Theorem˜1.3 all $S_{m}$ are finite and so define
$f_{m}:\mathcal{A}^{m}\to k^{|S_{m}|},\qquad f_{m}(\underline{a})_{i}=s_{i}(\underline{a}).$
Since $k$ is algebraically closed, it admits quantifier elimination. Conditions (i) and (ii) of Theorem˜2.1 are verified exactly as in Theorem˜2.2.
For (iii), note that since a finite group is discrete it is by definition reductive and has closed orbits. Hence by Theorem˜1.2, $k[\mathcal{A}^{m}]^{G}$ separates $G$-orbits. As before,
$f_{m}(\underline{a})=f_{m}(\underline{b})\iff\exists g\in G:\ g\underline{a}=\underline{b}.$
By Theorem˜1.1 and since $\text{Aut}(\mathcal{A})_{\mathcal{F}}=\text{Aut}(\mathcal{A})_{(\mathcal{F},E)}$, $\underline{a}$ and $\underline{b}$ have the same $\emptyset$-type in $(\mathcal{F},E)$. ∎
For an associative $k$-algebra $\mathcal{A}$, by an application of Wedderburn’s Theorem $\text{Aut}(\mathcal{A})$ is only finite in the case $\mathcal{A}\cong\mathbb{C}^{n}$. However finite subgroups of $\text{Aut}(\mathcal{A})$ exist and their invariant algebras, can be studied ([comp, fin]). For non-associative algebras, $\text{Aut}(\mathcal{A})$ can even itself be finite (and non-trivial). For example in [low], all semisimple, two dimensional $k$-algebras have been classified and their corresponding automorphism groups (some of which are finite) and invariant algebras calculated explicitly.
We finish with a statement using compact $\mathbb{R}$-rational points of linear algebraic groups without the necessity to extend the language by $\leq$. For simplicity we write it this time for $\mathbb{R}$ and $\mathbb{C}$. Generalization should again be possible, by translating everything into notions that can be transferred as first order logic statement.
Theorem 2.4.
Let $\mathcal{A}$ be a finite-dimensional $\mathbb{C}$-algebra. Assume $G<\mathrm{Aut}(\mathcal{A})_{\mathcal{F}}$ is a connected reductive linear algebraic group defined over $\mathbb{R}$ and let the $\mathbb{R}$-rational points $G_{\mathbb{R}}$ be compact. Let $S_{m}$ be a minimal generating set of $\mathbb{C}[\mathcal{A}^{m}]^{G}$ for each $m\in\mathbb{N}$, and let $(\mathcal{F},E)$ be an extended language such that the interpretation of each $e\in E$ in $\mathcal{A}$ lies in $\iota(\mathbb{C}[\mathcal{A}^{\text{arity}(e)}]^{G})$ and for all $s\in S_{m}$, $\iota(s)$ corresponds to evaluation of a $(\mathcal{F},E)$-term. Then the $(\mathcal{F},E)$-structure $\mathcal{A}$ admits quantifier elimination.
Proof.
By reductiveness of $G$ and Theorem˜1.3, each $S_{m}$ is finite, so define
$f_{m}:\mathcal{A}^{m}\to\mathbb{C}^{|S_{m}|},\qquad f_{m}(\underline{a})_{i}=s_{i}(\underline{a}).$
We want apply Theorem˜2.1 from the $(\mathcal{F},E)$ structure $\mathcal{A}$ to the $\mathcal{F}$ structure $\mathbb{C}$.
Condition (i) and (ii) can be shown exactly as in the previous proofs.
The only non trivial part again is (iii). By definition on $G$ Theorem˜1.2 implies:
$f_{m}(\underline{a})=f_{m}(\underline{b})\Rightarrow i(\underline{a})=i(\underline{b})\forall i\in\mathbb{C}[\mathcal{A}^{m}]^{G}\Leftrightarrow\overline{\mathrm{Orb}(\underline{a},G)}\cap\overline{\mathrm{Orb}(\underline{b},G)}\neq\emptyset$
So take a sequence $(g_{n})_{n\in\mathbb{N}}\subset G$, such that $g_{n}\cdot\underline{a}\to\underline{c}\in\overline{\mathrm{Orb}(\underline{b},G)}$. As $G$ is defined over $\mathbb{R}$ by the Hilbert-Mumford stability criterion (see [GIT] and [KN]) if $\underline{c}\notin\mathrm{Orb}(\underline{a},G)$ this sequence $g_{n}$ can be chosen from a real one parameter subgroup $H<G$. However, since $G$ is reductive, compactness of $G_{\mathbb{R}}$ implies that $G$ is $\mathbb{R}$-anisotropic i.e. contains no non-trivial $\mathbb{R}$-split tori $\mathbb{G}_{k}<G$ (see for example [borel]). Thus $H$ is trivial and therefore $\underline{c}\in\mathrm{Orb}(\underline{a},G)$. Repeating the argument for $\underline{b}$ we finally get the desired result:
$\exists g\in G:\ g\cdot\underline{a}=\underline{b}$
By Theorem˜1.1 this implies that $\underline{a}$ and $\underline{b}$ are of the same $\emptyset$-type in $\mathcal{F}$ and since all functions interpreted by $E$ are $G$-invariant, $\underline{a}$ and $\underline{b}$ are also of the same $\emptyset$-type in $(\mathcal{F},E)$. ∎
3 Acknowledgements
I want to greatly thank my supervisor Prof. Tim Netzer, for many helpful comments on a preliminary version of this article.
I also gratefully acknowledge funding by the Austrian Academy of Sciences (OAW), through a DOC-Scholarship.